The mean dying rate of mice administered yo a given dosage of an experimental drug is 45with a variance of 25. What is the probability randomly selected 55 mice, which received the drug, more than 50 will die? Assume normality

1
Expert's answer
2022-04-14T08:07:50-0400

Let X=X= the number of mice that died:

"X\u223cN(\u03bc,\u03c3 \n2\n )"

Given "\mu=45, \sigma^2=25"

"P(X>50)=1\u2212P(X\u226450)\n=1-P(Z\\leq \\dfrac{50-45}{\\sqrt{25}})=1-P(Z\\leq 1)=1\u2212P(Z\u2264 \n25\n\u200b\n \n50\u221245\n\u200b\n )=1\u2212P(Z\u22641)\n\\approx0.158655"